1. Find out the following values from the graph: profit, total costs, labor cost, capital cost. MC ATC -MR - AVC 30
Answer #1

Answer

The firm produces at MR=MC

Where

Q=30 units and P=$7

Profit=TR-TC

TC=ATC*Q=6*30=180

TR=7*30=210

Profit=210-180=$30

========

TC=ATC*Q=6*30=180

=====

Labour cost =variable cost =AVC*Q=5*30=$150

As labour is variable in the short run.

=====

Capital cost=fixed cost =TC-VC=180-150=$30

Capital is fixed in the short run